OpenStudy (anonymous):

m= 1 (0,0) find the equation of a given slope and a point write in y=mx+b form

OpenStudy (korosh23):

In here, you know m is the slope and o,o is the x and y value

OpenStudy (korosh23):

b stands for the place which the line intercepted the y axis

OpenStudy (korosh23):

Plug in the numbers inside the given equation and you will get

OpenStudy (korosh23):

y=1(units of x axis)+ 0

OpenStudy (anonymous):

y-y1=m(x-x1) substitute the points to x1 and y1

OpenStudy (anonymous):

b= -1 right

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y-(substitute y)= (substitute m) (x-(substitute x )

OpenStudy (anonymous):

what on earth?

OpenStudy (anonymous):

Y= 1(0) + -1

OpenStudy (anonymous):

the slope is one the point is \((0,0)\) the equation is \[y-0=1(x-0)\] which is only \(y=x\)

OpenStudy (anonymous):

how did you get y = x though?

OpenStudy (anonymous):

satellite already substituted the values

OpenStudy (anonymous):

youd you add zero to the other side

OpenStudy (anonymous):

would*

OpenStudy (anonymous):

y-0 = y

OpenStudy (anonymous):

no, the zeroes are given (0,0)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

^_^